OpenStudy (anonymous):

A group of 52 people attended a ball game. There was three times as many children as adults in the group. Set up a system of equations that represents the numbers of adults and children why attended the game and solve the system to find the number of childrem who were in the group.

OpenStudy (anonymous):

We can call a the number of adults and c the number of children so the equations would be a+c=52 3a=c any idea how to solve?

OpenStudy (anonymous):

solve for c right?

OpenStudy (anonymous):

Sorry to but in, but wouldn't the first equation be 3a+c=52?

OpenStudy (anonymous):

I mean a+3c=52 ?

OpenStudy (anonymous):

Thats the total number of people, adults+children=52

OpenStudy (anonymous):

Oh OK sorry

OpenStudy (anonymous):

It's ok, yes solve for c using substitution or elimination @tpenn

OpenStudy (anonymous):

I prefer elimination so much more.